Solve Triangle ABC. A=90, b=4, c=7. i figured out that the hypotenuse (leg a) is the square root of 65 by using the pythagorean theorem, but how do i find angle B and C?
Answer by stanbon(75887)

B=inverse tan (4/7)= 29.75 degrees
C= 190-29.75 = 60.25 degrees
Cheers,
Stan H.